OpenStudy (anonymous):

The pair of points is on the graph of an inverse variation. Find the missing value. (3, 7) and (x, 6)

hartnn (hartnn):

for the inverse variation xy =k = constant since 3,7 lies on it, 3*7 = k so, k =21 also since x,6 lies on xy = k x*6 = 21 get the value of x from here :)

hartnn (hartnn):

oh and @fredypaliku \(\large \frak Welcome ~ to ~ OpenStudy !!\)

OpenStudy (mathmale):

hartnn is right on track, and his final equation will produce the desired value (x value). xy = k is also correct as a way of writing a function involving inverse variation; writing y=k/x may be easier for some to recognize and understand.